# Informed Initialization for Bayesian Optimization and Active Learning

### v1

@@ 1 Introduction
Bayesian Optimization (BO) (Mockus et al., 1978; Jones et al., 1998; Frazier, 2018; Garnett, 2023) is a principled framework for sample-efficient global optimization of black-box functions with applications across diverse fields such as biological discovery (Griffiths and Hernández-Lobato, 2020; Stanton et al., 2022), materials science (Frazier and Wang, 2016; Attia et al., 2020; Ament et al., 2023b), online A/B testing (Agarwal et al., 2018; Letham et al., 2019; Feng et al., 2025), and machine learning hyperparameter optimization (HPO) (Snoek et al., 2014; Feurer et al., 2015). BO combines a probabilistic surrogate model—commonly a Gaussian Process (GP)—with an acquisition function to select where to evaluate the unknown objective function. In many applications, the runtime of a single black-box function evaluation may restrict the experimenter to a small number of batches – the number of sequential rounds of experiments – but many real-world experimental setups permit conducting multiple experiment simultaneously (e.g., on a parallel compute cluster, in a randomized controlled trial, or batch-testing multiple specimens in a lab).
The success of Bayesian Optimization in practice is highly sensitive to the quality of the surrogate model (Eriksson and Jankowiak, 2021; Hvarfner et al., 2023; Foldager et al., 2023). This is a challenge particularly during the early stages of optimization when few observations are available. In these early stages, a small set of inputs is typically selected at random, or via space-filling or quasi-random sampling strategies such as Latin Hypercube Sampling (LHS) or scrambled Sobol’ sequences (Bossek et al., 2020; Owen, 2023). While such strategies aim to achieve broad coverage of the input space for initialization of the surrogate model, they may not necessarily result in improving predictive accuracy, or reduce predictive variance, across the entire input space Ren and Sweet (2024). Moreover, they neglect a a second crucial aspect of modeling: the need to accurately infer the model’s hyperparameters (Zhang et al., 2019), such as the kernel lengthscales of a GP. With accurate hyperparameter estimates, variation in unimportant dimensions will have less influence on the selection of points in subsequent iterations of BO, leading to more sample-efficient optimization (Eriksson and Jankowiak, 2021; Hvarfner et al., 2023; Müller et al., 2023). Conversely, poor hyperparameter estimation may cause subsequent BO iterations to fail to make meaningful progress (Berkenkamp et al., 2019), which may be caused by misidentifying signal for noise, exploring irrelevant dimensions, or returning poor terminal recommendations (Hvarfner et al., 2022).
@@ 2.1 Gaussian Processes
To model differences in variable importance in GPs with stationary kernels, each input dimension is commonly scaled by a lengthscale hyperparameter $\ell_{i}$, a practice known as Automatic Relevance Determination (ARD) (Williams and Rasmussen, 1995). Additional, optional hyperparameters include a learnable noise variance $\sigma_{\varepsilon}^{2}$ and signal variance $\sigma_{f}^{2}$. The full set of hyperparameters $\bm{\theta}=\{\bm{\ell},\sigma_{\varepsilon}^{2},\sigma_{f}^{2}\}$ may be learned either by maximizing the marginal likelihood $p(\mathcal{D}\mid\bm{\theta})$ (Maximum Likelihood Estimation, MLE) or by incorporating hyperpriors $p(\bm{\theta})$ to perform Maximum A Posteriori (MAP) estimation. Alternatively, a fully Bayesian treatment (Osborne, 2010; Lalchand and Rasmussen, 2020) integrates over $\bm{\theta}$ to approximate the full Bayesian posterior distribution using Markov Chain Monte Carlo (MCMC) methods, thereby explicitly accounting for hyperparameter uncertainty. For additional background on GPs, see (Rasmussen and Williams, 2005).
@@ 2.2 Bayesian Optimization
Bayesian Optimization (BO) (Frazier, 2018) is a sample-efficient framework for finding the maximizer $\bm{x}_{*}=\mathop{\mathrm{arg\,max}}_{\bm{x}\in\mathcal{X}}f(\bm{x})$ of a black-box function $f:\mathcal{X}\rightarrow\mathbb{R}$ over a $D$-dimensional input space $\mathcal{X}=[0,1]^{D}$. The function $f$ is assumed to be expensive to evaluate and observable only through noisy point-wise measurements, $y(\bm{x})=f(\bm{x})+\varepsilon$, where $\varepsilon\sim\mathcal{N}(0,\sigma_{\varepsilon}^{2})$.
At the core of BO is an acquisition function, which uses a surrogate model to quantify the (expected) utility of candidate points $\bm{x}$. Acquisition functions balance exploration and exploitation typically through greedy heuristics. Popular examples include Expected Improvement (EI) (Jones et al., 1998; Bull, 2011) and its numerically stable variant, LogEI (Ament et al., 2023a), as well as the Upper Confidence Bound (UCB) (Srinivas et al., 2012, 2010). In batch BO, multiple points are selected in parallel to accelerate data collection. This is often achieved by computing and jointly optimizing a (quasi-)MC estimate of the utility $u$ associated with acquisition function over the full batch ${\bm{X}}=\{\bm{x}_{1},\bm{x}_{2},...,\bm{x}_{q}\}$ of size $q$ (Wilson et al., 2017, 2020; Balandat et al., 2020).
@@ 2.3 Bayesian Active Learning
Bayesian Active Learning (BAL) and Bayesian Experimental Design (BED) (Chaloner and Verdinelli, 1995) aims to improve predictive models by strategically selecting data points that are most informative, either with regard to the model or to future predictions. A central quantity is the Expected Information Gain (EIG):
@@
Bayesian Active Learning by Disagreement (BALD) (Houlsby et al., 2011; Kirsch et al., 2019) selects query points that maximize the mutual information between model predictions and hyperparameters $\bm{\theta}$:
@@
BALD identifies locations where models within an ensemble exhibit the greatest disagreement in predictive uncertainty. In the GP setting, this often leads to axis-aligned queries when there is high uncertainty in the lengthscales, and to repeated queries when observation noise is highly uncertain. Notably, BALD is model-agnostic and can be applied to a wide range of surrogate models and hyperparameters, including subspace models (Garnett et al., 2014) and additive decompositions (Hvarfner et al., 2023; Gardner et al., 2017).
Negative Integrated Posterior Variance and Expected Predictive Information Gain
The Negative Integrated Posterior Variance (NIPV) (Chen and Zhou, 2014; Seo et al., 2000) criterion selects queries that minimize the expected posterior variance over a test distribution $p_{*}(\bm{x})$:
@@
Similarly, Expected Predictive Information Gain (EPIG) (Bickford Smith et al., 2023) selects queries that minimize the expected predictive entropy over $p_{*}(\bm{x})$:
$\text{EPIG}(\bm{x};p_{*})=-\mathbb{E}_{\bm{x}_{*}\sim p_{*}}\left[\text{H}[y(\bm{x})]\;|\;\bm{x}_{*},\mathcal{D}\right].$ (4)
Both objectives promote the selection of data that reduces uncertainty over the test distribution, but without considering the effect on hyperparameter learning. The test distribution encodes how much emphasis is put on different parts of the domain. It can be specified by subject matter experts; in the case of no prior knowledge it is typically the uniform distribution. Throughout the remainder of the paper, we will exclusively consider NIPV with a uniform $p_{*}$.
@@ 3 Related Work
Initialization has received surprisingly little attention in the context of BO. Bossek et al. (2020) conducts a study on the effect of various random initial designs on BO performance. Alternatively, minimax or maximin criteria (Johnson et al., 1990) may be used to accomplish evenly distributed designs. Maybe closest to our work is (Zhang et al., 2019), which proposes LHS-Beta, an initial design criterion which alters samples drawn by LHS to achieve pairwise distances between points which matches a Beta distribution. LHS-Beta pursues diverse pairwise distances in the data, in order to best learn the lengthscale of a GP with an isotropic kernel. Zimmerman (2006); Müller and Zimmerman (1999) address the problem of learning parameters of Kriging estimators using a empirical estimates of optimal experimental design criteria, limiting candidates to a fixed grid of points.
Entropy-maximizing (Guestrin et al., 2005; MacKay, 1992, 1995) or variance-minimizing (Park et al., 2024) designs have been explored in active learning for optimal sensor placement (Krause et al., 2006, 2008) and other applications involving GPs, such as geostatistics (Sauer et al., 2023) and contour finding (Cole et al., 2023). Moreover, parameter-related EIG criteria are a bedrock of the broader topic of BED (Bickford Smith et al., 2023; Rainforth et al., 2024; Kirsch et al., 2021; Chaloner and Verdinelli, 1995), which focuses on selecting data that is most informative about model parameters or future predictions. These prediction or model-oriented criteria have yet to see widespread use in BO, particularly for initialization. However, information-theoretic acquisition functions (Hennig and Schuler, 2012; Hernández-Lobato et al., 2014; Wang and Jegelka, 2017; Moss et al., 2021; Hvarfner et al., 2022; Tu et al., 2022; Neiswanger et al., 2021, 2024) address the optimization problem from an information theoretic perspective, albeit not with a primary focus on initialization or model predictive performance.
The problem of actively learning model hyperparameters during BO (post initialization) has previously been investigated by Hvarfner et al. (2023), who propose a combined BO-BAL framework to actively learn the hyperparameters of the GP along with the optimum, and demonstrate that better-calibrated surrogates significantly enhance BO performance. Houlsby et al. (2011) proposes BALD, an active learning acquisition functions for hyperparameters in preference learning in GPs. Riis et al. (2022) proposes a a Query-By-Committee-oriented acquisition function for BAL in GPs. Lastly, Berkenkamp et al. (2019); Ziomek et al. (2024) address BO performance under hyperparameter uncertainty from a theoretical perspective, proving regret bounds when hyperparameters of the objective are unknown.
@@ 4 Method
We consider the case of large-batch, few-shot BO, where the batch size $q$ is large ($q\geq 8$), and only a small number of sequential batches $B$ can be evaluated (often only one for initialization and one for BO, so $B=2$). With only a handful of batches, there is little opportunity to correct for poor initialization. If the initial design fails to reduce uncertainty in key hyperparameters or leaves large regions of the input space unexplored, both hyperparameter inference and optimization become challenging. This setting is common if evaluations are lengthy but can be effectively parallelized, which is the case for instance in online A/B tests or lab experiments for biology Stanton et al. (2022); González-Duque et al. (2024) or material design Yik et al. (2025).
@@
While Sobol and other space-filling sequences are designed to uniformly cover the input space, this property alone does not imply that the resulting model will be desirably helpful for subsequent BO. In fact, space-filling is not synonymous with informativeness: a model trained on a space-filling design may still exhibit high predictive uncertainty in regions most relevant for optimization. For example, random designs increasingly sample near the boundaries of the search space as dimensionality grows Swersky (2017); Köppen (2000), whereas a point located at the center of the search space may be far more informative for optimization, as it can yield a greater reduction in predictive variance throughout the search space.
To this point, a strategy that explicitly reduces predictive uncertainty within the region of interest Bickford Smith et al. (2023)—namely, the search space—not only seems better suited for our goals, but is also better aligned with both theoretical (Srinivas et al., 2010; Berkenkamp et al., 2019) and information gain-based criteria, as well as the practical objective of achieving a more informed, in an informal sense, model post-initialization Garnett (2023). Methods like NIPV (Zhang et al., 2019) which minimize predictive variance over the input space achieve this goal better than Sobol, and yields designs that are spread out but not necessarily informative for model calibration and hyperparameter learning.
On the other hand, methods that focus exclusively on hyperparameter informativeness, such as BALD (Houlsby et al., 2011), tend to cluster queries along specific axes in order to resolve lengthscales, thereby sacrificing coverage and potentially overlooking important regions of the search space. When additional hyperparameters are present, such as a learnable noise variance, BALD may even select duplicate queries at the same location to better estimate said noise variance. While such behavior is not inherently problematic, it can be undesirable in the few-shot setting, where initialization must address hyperparameter learning and space-filling simultaneously.

### v2

@@ 1 Introduction
Bayesian Optimization (BO) (Mockus1978; jonesei; frazier2018tutorial; garnett_bayesoptbook_2023) is a principled framework for sample-efficient global optimization of black-box functions with applications across diverse fields such as biological discovery (griffiths2020constrained; pmlr-v162-stanton22a), materials science (frazier2016bayesian; attia2020closed; ament2023sustainableconcretebayesianoptimization), online A/B testing (agarwal2018online; letham2019constrained; feng2025fastandslow), and machine learning hyperparameter optimization (HPO) (pmlr-v32-snoek14; Feurer2015InitializingBH). BO combines a probabilistic surrogate model—commonly a Gaussian Process (GP)—with an acquisition function to select where to evaluate the unknown objective function. In many applications, the runtime of a single black-box function evaluation may restrict the experimenter to a small number of batches – the number of sequential rounds of experiments – but many real-world experimental setups permit conducting multiple experiment simultaneously (e.g., on a parallel compute cluster, in a randomized controlled trial, or batch-testing multiple specimens in a lab).
The success of Bayesian Optimization in practice is highly sensitive to the quality of the surrogate model (eriksson2021high; hvarfner2023selfcorrecting; pmlr-v216-foldager23a). This is a challenge particularly during the early stages of optimization when few observations are available. In these early stages, a small set of inputs is typically selected at random, or via space-filling or quasi-random sampling strategies such as Latin Hypercube Sampling (LHS) or scrambled Sobol’ sequences (10.1145/3377930.3390155; owen2023practicalqmc). While such strategies aim to achieve broad coverage of the input space for initialization of the surrogate model, they may not necessarily result in improving predictive accuracy, or reduce predictive variance, across the entire input space ren2024optimalinitializationbatchbayesian. Moreover, they neglect a a second crucial aspect of modeling: the need to accurately infer the model’s hyperparameters (zhang2021ddd), such as the kernel lengthscales of a GP. With accurate hyperparameter estimates, variation in unimportant dimensions will have less influence on the selection of points in subsequent iterations of BO, leading to more sample-efficient optimization (eriksson2021high; hvarfner2023selfcorrecting; pmlr-v202-muller23a). Conversely, poor hyperparameter estimation may cause subsequent BO iterations to fail to make meaningful progress (berkenkamp2019noregret), which may be caused by misidentifying signal for noise, exploring irrelevant dimensions, or returning poor terminal recommendations (hvarfner2022joint).
@@ 2.1 Gaussian Processes
To model differences in variable importance in GPs with stationary kernels, each input dimension is commonly scaled by a lengthscale hyperparameter $\ell_{i}$, a practice known as Automatic Relevance Determination (ARD) (NIPS1995_7cce53cf). Additional, optional hyperparameters include a learnable noise variance $\sigma_{\varepsilon}^{2}$ and signal variance $\sigma_{f}^{2}$. The full set of hyperparameters $\bm{\theta}=\{\bm{\ell},\sigma_{\varepsilon}^{2},\sigma_{f}^{2}\}$ may be learned either by maximizing the marginal likelihood $p(\mathcal{D}\mid\bm{\theta})$ (Maximum Likelihood Estimation, MLE) or by incorporating hyperpriors $p(\bm{\theta})$ to perform Maximum A Posteriori (MAP) estimation. Alternatively, a fully Bayesian treatment (osborne2010bayesian; lalchand2020approximate) integrates over $\bm{\theta}$ to approximate the full Bayesian posterior distribution using Markov Chain Monte Carlo (MCMC) methods, thereby explicitly accounting for hyperparameter uncertainty. For additional background on GPs, see (10.5555/1162254).
@@ 2.2 Bayesian Optimization
Bayesian Optimization (BO) (frazier2018tutorial) is a sample-efficient framework for finding the maximizer $\bm{x}_{*}=\mathop{\mathrm{arg\,max}}_{\bm{x}\in\mathcal{X}}f(\bm{x})$ of a black-box function $f:\mathcal{X}\rightarrow\mathbb{R}$ over a $D$-dimensional input space $\mathcal{X}=[0,1]^{D}$. The function $f$ is assumed to be expensive to evaluate and observable only through noisy point-wise measurements, $y(\bm{x})=f(\bm{x})+\varepsilon$, where $\varepsilon\sim\mathcal{N}(0,\sigma_{\varepsilon}^{2})$.
At the core of BO is an acquisition function, which uses a surrogate model to quantify the (expected) utility of candidate points $\bm{x}$. Acquisition functions balance exploration and exploitation typically through greedy heuristics. Popular examples include Expected Improvement (EI) (jonesei; JMLR:v12:bull11a) and its numerically stable variant, LogEI (ament2023unexpected), as well as the Upper Confidence Bound (UCB) (Srinivas_2012; srinivas2010ucb). In batch BO, multiple points are selected in parallel to accelerate data collection. This is often achieved by computing and jointly optimizing a (quasi-)MC estimate of the utility $u$ associated with acquisition function over the full batch ${\bm{X}}=\{\bm{x}_{1},\bm{x}_{2},...,\bm{x}_{q}\}$ of size $q$ (wilson2017reparam; wilson2020efficiently; balandat2020botorch).
@@ 2.3 Bayesian Active Learning
Bayesian Active Learning (BAL) and Bayesian Experimental Design (BED) (chaloner1995bed) aim to improve predictive models by selecting data points that are most informative, either with regard to the model or to future predictions. A key quantity is the Expected Information Gain (EIG):
@@
Bayesian Active Learning by Disagreement (BALD) (houlsby2011bald; kirsch2019batchbald) selects query points that maximize the mutual information between model predictions and hyperparameters $\bm{\theta}$:
@@
BALD identifies locations where models within an ensemble exhibit the greatest disagreement in predictive uncertainty. In the GP setting, this often leads to axis-aligned queries when there is high uncertainty in the lengthscales, and to repeated queries when observation noise is highly uncertain. Notably, BALD is model-agnostic and can be applied to a wide range of surrogate models and hyperparameters, including subspace models (garnett2014active) and additive decompositions (hvarfner2023selfcorrecting; pmlr-v54-gardner17a).
Negative Integrated Posterior Variance
The Negative Integrated Posterior Variance (NIPV) (sambu2000GPRactiveselection) criterion selects queries that minimize the expected posterior variance over a test distribution $p_{*}(\bm{x})$:
@@
Expected Predictive Information Gain
The Expected Predictive Information Gain (EPIG) (pmlr-v206-bickfordsmith23a) selects a candidate point $\bm{x}$ by maximizing the mutual information between its label $y(\bm{x})$ and the label $y(\bm{x}^{*})$ at a randomly drawn test input $\bm{x}_{*}\sim p_{*}(\bm{x})$,
$\mathrm{EPIG}(\bm{x};p_{*})=\mathbb{E}_{\bm{x}_{*}\sim p_{*}}\!\left[I\!\big(y(\bm{x});y(\bm{x}_{*})\mid\mathcal{D}\big)\right].$ (4)
which can equivalently be expressed in terms of an expectation over a difference of entropies. In its general, model-agnostic form, this constitutes of a nested expectation over both test inputs $\bm{x}_{*}$ and outcomes and selected locations $y(\bm{x})$:
$\mathrm{EPIG}(\bm{x};p_{*})=\mathbb{E}_{\bm{x}_{*}\sim p_{*}}\!\left[\text{H}\!\left[y(\bm{x}_{*})\mid\mathcal{D}\right]-\mathbb{E}_{y(\bm{x})\mid\mathcal{D}}\!\left[\text{H}\!\left[y(\bm{x}_{*})\mid\mathcal{D},\,\bm{x},\,y(\bm{x})\right]\right]\right].$ (5)
EPIG for Gaussian Processes
Standard GP regression with Gaussian likelihoods constitutes a special case for EPIG, as the predictive variance (and therefore the predictive entropy) after conditioning on $(\bm{x},y(\bm{x}))$ is independent of the realized value $y(\bm{x})$. Thus,
$\text{H}\!\left[y(\bm{x}_{*})\mid\mathcal{D},\,\bm{x},\,y(\bm{x})\right]=\text{H}\!\left[y(\bm{x}_{*})\mid\mathcal{D},\,\bm{x}\right],$
and the inner expectation over $y(\bm{x})$ in (5) disappears. EPIG therefore simplifies to an expected reduction in predictive entropy:
$\mathrm{EPIG}_{\mathrm{GP}}(\bm{x};p_{*})=\mathbb{E}_{\bm{x}_{*}\sim p_{*}}\!\left[\text{H}\!\left[y(\bm{x}_{*})\mid\mathcal{D}\right]-\text{H}\!\left[y(\bm{x}_{*})\mid\mathcal{D},\,\bm{x}\right]\right].$ (6)
We will use this GP-specific form of EPIG throughout the remainder of this paper.
Both NIPV and EPIG promote the selection of data that reduces uncertainty over the test distribution, but without considering the effect on hyperparameter learning. The test distribution encodes how much emphasis is put on different parts of the domain. It can be specified by subject matter experts; in the case of no prior knowledge it is typically the uniform distribution. Throughout the remainder of the paper, we will exclusively consider NIPV with a uniform $p_{*}$.
@@ 3 Related Work
Initialization has received surprisingly little attention in the context of BO. 10.1145/3377930.3390155 conducts a study on the effect of various random initial designs on BO performance. Alternatively, minimax or maximin criteria (JOHNSON1990131) may be used to accomplish evenly distributed designs. Maybe closest to our work is (zhang2021ddd), which proposes LHS-Beta, an initial design criterion which alters samples drawn by LHS to achieve pairwise distances between points which matches a Beta distribution. LHS-Beta pursues diverse pairwise distances in the data, in order to best learn the lengthscale of a GP with an isotropic kernel. zimmerman2006optimal; muller1999optimal address the problem of learning parameters of Kriging estimators using a empirical estimates of optimal experimental design criteria, limiting candidates to a fixed grid of points.
Entropy-maximizing (guestrin2005near; mackay1992information; MACKAY199573) or variance-minimizing (park2024activelearningpiecewisegaussian) designs have been explored in active learning for optimal sensor placement (krause2006near; JMLR:v9:krause08a) and other applications involving GPs, such as geostatistics (Sauer02012023) and contour finding (Cole01012023). Moreover, parameter-related EIG criteria are a bedrock of the broader topic of BED (pmlr-v206-bickfordsmith23a; rainforth2024modern; kirsch2021jepig; chaloner1995bed), which focuses on selecting data that is most informative about model parameters or future predictions. These prediction or model-oriented criteria have yet to see widespread use in BO, particularly for initialization. However, information-theoretic acquisition functions (entropysearch; pes; wang2017maxvalue; moss2021gibbon; hvarfner2022joint; tu2022joint; pmlr-v139-neiswanger21a; 10.5555/3600270.3601798) address the optimization problem from an information theoretic perspective, albeit not with a primary focus on initialization or model predictive performance.
The problem of actively learning model hyperparameters during BO (post initialization) has previously been investigated by hvarfner2023selfcorrecting, who propose a combined BO-BAL framework to actively learn the hyperparameters of the GP along with the optimum, and demonstrate that better-calibrated surrogates significantly enhance BO performance. houlsby2011bald proposes BALD, an active learning acquisition functions for hyperparameters in preference learning in GPs. riis2022bayesian proposes a a Query-By-Committee-oriented acquisition function for BAL in GPs. Lastly, berkenkamp2019noregret; NEURIPS2024_9cf904c8 address BO performance under hyperparameter uncertainty from a theoretical perspective, proving regret bounds when hyperparameters of the objective are unknown.
@@ 4 Method
We consider the case of large-batch, few-shot BO, where the batch size $q$ is large ($q\geq 8$), and only a small number of sequential batches $B$ can be evaluated (often only one for initialization and one for BO, so $B=2$). With only a handful of batches, there is little opportunity to correct for poor initialization. If the initial design fails to reduce uncertainty in key hyperparameters or leaves large regions of the input space unexplored, both hyperparameter inference and optimization become challenging. This setting is common if evaluations are lengthy but can be effectively parallelized, which is the case for instance in online A/B tests or lab experiments for biology pmlr-v162-stanton22a; gonzalez2024survey or material design yik2025battery.
@@
While Sobol and other space-filling sequences are designed to uniformly cover the input space, this property alone does not imply that the resulting model will be desirably helpful for subsequent BO. In fact, space-filling is not synonymous with informativeness: a model trained on a space-filling design may still exhibit high predictive uncertainty in regions most relevant for optimization. For example, random designs increasingly sample near the boundaries of the search space as dimensionality grows swersky2017improving; koppencurse, whereas a point located at the center of the search space may be far more informative for optimization, as it can yield a greater reduction in predictive variance throughout the search space.
To this point, a strategy that explicitly reduces predictive uncertainty within the region of interest pmlr-v206-bickfordsmith23a—namely, the search space—not only seems better suited for our goals, but is also better aligned with both theoretical (srinivas2010ucb; berkenkamp2019noregret) and information gain-based criteria, as well as the practical objective of achieving a more informed, in an informal sense, model post-initialization garnett_bayesoptbook_2023. Methods like NIPV (zhang2021ddd) which minimize predictive variance over the input space achieve this goal better than Sobol, and yields designs that are spread out but not necessarily informative for model calibration and hyperparameter learning.
On the other hand, methods that focus exclusively on hyperparameter informativeness, such as BALD (houlsby2011bald), tend to cluster queries along specific axes in order to resolve lengthscales, thereby sacrificing coverage and potentially overlooking important regions of the search space. When additional hyperparameters are present, such as a learnable noise variance, BALD may even select duplicate queries at the same location to better estimate said noise variance. While such behavior is not inherently problematic, it can be undesirable in the few-shot setting, where initialization must address hyperparameter learning and space-filling simultaneously.